Recent Log Entries

MQA3A29 Trim Card Swapped

Log Number: 3814776Logbook(s): Entry Makers: moserBody: 

Magnet setpoint was stuck. Not taking demand, not cycling. Swapped trim card, all is well.

1L25-8 Vacuum Degredation at ODVH=20.5Mv/m

Log Number: 3814775Logbook(s): Entry Makers: kernsBody: 

Crew chief and I played with trying to keep this zone from dropping HV . Last week I found that cavity 8 would drop all eight out at anything above 10Mv/m. Gradually lowered from 20.5 Mv/m until gradient stopped affecting the vacuum with all the others at ODVH. Didn't change ODVH due to forgetfullnes and Lem put it back at 20.5 and it dropped out several dozen times until MCC made not of it. D. Moser will see how it holds up at 10Mv/m and we'll take it in the morning !!!

CEBAF Magnets from 3/24 Downloaded, Cycled

Log Number: 3814773Logbook(s): Entry Makers: moserBody: 

Used all save 21402 per the COVID restart plan, ATLis 20836.

First, selected "All Magnets"; grep -v 0[LIR]; grep BDL; Get Current Values. Download.... Performed a back-up all-save (in addition to auto-save before download). #21431

Next, cycled all box supplies. Then trims.

HYST AREA 6-2-1 Cycles Magnets

Log Number: 3814771Logbook(s): Body: 
HYST AREA Version 6-2-1 has FAILED to cycle:

MQD1L05 -- NL01B02 -- nl6mag -- Failed to set after 2 cycle attempts

MQA3A29 -- SA01B06 -- ea6mag -- Failed to set after 2 cycle attempts

MBDAS10H -- W102B10 -- wa1mag -- Failed to set after 2 cycle attempts

MQA4A12 -- W301B13 -- wa3mag -- Failed to set after 2 cycle attempts

MQA4A21 -- W501B11 -- wa5mag -- Failed to set after 2 cycle attempts
MQA6A21 -- W501B11 -- wa5mag -- Failed to set after 2 cycle attempts

Box Supplies Were Cycled

Log Number: 3814769Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.297 A/s
ARC2A Ramp Rate = 6.895 A/s
ARC3A Ramp Rate = 9.835 A/s
ARC4A Ramp Rate = 11.047 A/s
ARC5A Ramp Rate = 10.161 A/s
ARC6A Ramp Rate = 9.135 A/s
ARC7A Ramp Rate = 10.203 A/s
ARC8A Ramp Rate = 11.950 A/s
ARC9A Ramp Rate = 8.959 A/s
ARCAA Ramp Rate = 12.244 A/s
LAM1C Ramp Rate = 10.500 A/s
LAM3C Ramp Rate = 11.854 A/s
MYA2T01 Ramp Rate N/A
MYA4T01 Ramp Rate N/A
MYA6T01 Ramp Rate N/A
MYAAT01 Ramp Rate N/A
RSEP8A Ramp Rate = 10.471 A/s
RSEP9A Ramp Rate = 9.366 A/s

Autolog restore

Log Number: 3814768Logbook(s): Tags: Body: saver ran from opsl632
Data from: /usr/opdata/burt/allsnap/03-24-20_07:04:35
Restored file: /tmp/07-14-20_21:24:20.snap


3/24 CEBAF magnet settings

Autolog save

Log Number: 3814767Logbook(s): Tags: Body: saver ran from opsl632

Saved: moser 07-14-20 21:24:07 {A:1 B:5 C:5 E:1050 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in. Autosave prior to restore.

Autolog save

Log Number: 3814766Logbook(s): Tags: Body: saver ran from opsl632

Saved: moser 07-14-20 21:23:08 {A:1 B:5 C:5 E:1050 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

LEM North Linac - failure

Log Number: 3814764Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Tue Jul 14 20:43:44 2020. The operation terminated after 0:00:41 with the following results.
Cavity summarySuccess: 186 Abort: 14 Fail: 0

Cycling the (CEBAF) world

Log Number: 3814763Logbook(s): Entry Makers: moserBody: 

All available boxes first -- Figure 1. Then trims. Figure 2.

Figure 3: had six trims that didn't cycle. 3 were successful cycling outside of hystarea, the other three need new trim cards.

Autolog restore

Log Number: 3814761Logbook(s): Tags: Body: saver ran from opsl632
Data from: /usr/opdata/burt/allsnap/07-14-20_19:44:01
Restored file: /tmp/07-14-20_19:51:56.snap


Autolog save

Log Number: 3814760Logbook(s): Tags: Body: saver ran from opsl632

Saved: kazimi 07-14-20 19:44:01 {A:1 B:5 C:5 E:1050 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

BLA HCO update

Log Number: 3814759Logbook(s): Body: 

BLA HCO Update:

I ran confidence tests on Halls A, B, and D. Nullifying the calibration coefficient, the test source current readbacks were 5.12uA, 5006 nA, and 4936 nA, respectively. This is sufficient to check the cavities off in HCO.

I also verified the tune on Hall B 2C24, because of the vacuum work done there. Without adjusting anything, the cavity was tuned to 1497.223 MHz, which is close enough, so I left it alone.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3814758Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: eforman
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C0

LEM North Linac - failure

Log Number: 3814757Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Tue Jul 14 19:38:18 2020. The operation terminated after 0:00:29 with the following results.
Cavity summarySuccess: 186 Abort: 14 Fail: 0

HYST AREA 6-2-1 Cycles Magnets

Log Number: 3814756Logbook(s): Body: 
HYST AREA Version 6-2-1 has FAILED to cycle:

MQD0R03 -- NA01B10 -- wa6mag -- Failed to set after 2 cycle attempts

MQD0R09 -- NL01B01 -- nl6mag -- Failed to set after 2 cycle attempts

MQD1L05 -- NL01B02 -- nl6mag -- Failed to set after 2 cycle attempts

MQD9S02 -- NL27B21 -- nl3mag -- Failed to set after 2 cycle attempts

MQA3A29 -- SA01B06 -- ea6mag -- Failed to set after 2 cycle attempts

MQA4A12 -- W301B13 -- wa3mag -- Failed to set after 2 cycle attempts

Pages